List the first four terms of the sequence defined by an =(n + 1)!/2n for for n is greater than or equal to 1.
Answer by stanbon(75887)

List the first four terms of the sequence defined by an =(n + 1)!/2n for for n is greater than or equal to 1.
-----
a(1) = (1+1)!/(2*1) = 2/2 = 1
a(2) = (2+1)!/(2*2) = 6/4 = 3/2
a(3) = (3+1)!/(2*3) = 24/6 = 4
a(4) = (4+1)!/(2*4) = 120/8 = 15
============================
Cheers,
Stan H.
====================